OpenStudy (anonymous):

a 68g gold bar is cut into 3 equal pieces how does the density change

OpenStudy (gorv):

density = mass/volume

OpenStudy (gorv):

new mass= m/3

OpenStudy (gorv):

new volume = v/3

OpenStudy (gorv):

new density=(m/3)/(v/3)=m/v so density remain same

OpenStudy (anonymous):

okay thanks! now i get it

OpenStudy (gorv):

welcome :)